Chapter 03, Problem 021 An ant, crazed by the Sun on a hot Texas afternoon, darts over an xy plane scratched in the dirt. The x and y components of four consecutive darts are the folowing, au in centimeters: (25.0, 47.0). (br.-64.0), (-19.0. Су), (-81.0,-53.0) The overall displacement of the four darts has the xy components (-140, 32.0).What are (a) b and (b) cy? What are the (e) magnitude and (d) angle (counterclockwise from the positive direction of the x axis) of the overall displacement?
Answer #1

overrightarrow{r}=overrightarrow{r_{1}}+overrightarrow{r_{2}}+overrightarrow{r_{3}}+overrightarrow{r_{4}}

a)

25+bx-19-81=-140

bx=-65 cm

b)

47-64+cy-53=-32

cy=38 cm

c)

r=sqrt[(-65)2+382]=75.3 cm

d)

)-180-30.3-149.7 e-tan-1 θ = tan -180-30.3 149.70 -65 Counterclockwise from +x-axis